# Generating 6-D Trajectories for Omnidirectional Multirotor Aerial Vehicles in Cluttered Environments

### v1

@@ III-D Rotation Parameterization
where $\mathbf{U}:=\begin{bmatrix}-\mathbf{r}&w\mathbf{I}+\mathbf{r}^{\wedge}\end{bmatrix}\in\mathbb{R}^{3\times 4}$ and $\mathbf{G}:=\begin{bmatrix}\frac{\partial w}{\partial\bm{\sigma}}&\frac{\partial x}{\partial\bm{\sigma}}&\frac{\partial y}{\partial\bm{\sigma}}&\frac{\partial z}{\partial\bm{\sigma}}\end{bmatrix}\in\mathbb{R}^{3\times 4}$. Then, we can calculate the rotation-related penalty terms.
@@ IV-B Parameter Settings
To achieve whole-body obstacle avoidance, as Section III-A mentions, we approximate the vehicle’s shape as a convex polyhedron that envelops the body. Here we approximate it as a cuboid (Fig. 5a) whose dimensions are $l_{x}=l_{y}=1.1\text{m},l_{z}=0.42\text{m}$, according to the size of OmniHex. The coordinates of its eight vertices in $\mathcal{F}_{b}$ are $\tilde{\bm{v}}_{l}=\begin{bmatrix}\pm\frac{l_{x}}{2}&\pm\frac{l_{y}}{2}&\pm\frac{l_{z}}{2}\end{bmatrix}^{\top}$. We set $s=4$ to ensure the smoothness of control inputs according to (8). The initial value of $\mathbf{q}^{\sigma}$ of trajectory optimization is set to $\mathbf{0}$. The other optimization settings for simulations are listed in Table I.
Figure 7: Trajectory generation results. (a) shows a trajectory generated in Scenario B (from the top view), and (b) shows that in Scenario C (from the side view).
@@ IV-C Simulation Results
In this section, simulations are carried out in 3 virtual scenarios to test the proposed method’s effectiveness and efficiency. The three virtual scenarios are described as follows.
Scenario A: We expect the vehicle to fly through a narrow, straight, vertical passage. The passage is 0.7m wide, which is narrower than the vehicle’s width ($l_{x}$ and $l_{y}$) and wider than its thickness ($l_{z}$). Therefore, the vehicle must tilt at a large angle to pass through without collision.
@@
Scenario C: We expect the vehicle to pass through a $30\text{m}\times 10\text{m}\times 8\text{m}$ area containing randomly distributed floating cuboid obstacles.
@@
TABLE II: Efficiency Evaluation
@@
C 18 23 37.0m 49ms 152ms 48 603ms
Fig. 6a shows the 6-D trajectory generated in Scenario A and the corresponding kinodynamic properties under the above settings. The generated trajectory allows the vehicle to smoothly and appropriately change its attitude to adapt to the narrow space in the passage, avoiding collision while flying toward the target. Fig. 6b presents the constrained kinodynamic properties. Although the constraints are relaxed, they are effectively satisfied in the resulting trajectory. Moreover, the velocity norm reaches $v_{\text{max}}$ most of the time, which shows that our method can fully exploit the performance of OMAVs. Note that due to the softening step, the resulting trajectory may slightly exceed the constraints [10], so it is better to reserve a certain margin in practice.
@@
For efficiency evaluation, we list the computation times of the results in Fig. 6 and 7 in Table II. $D$ is the straight-line distance between a trajectory’s start and end positions. $t_{\text{rrt}}$ is the time it takes to find an initial feasible path using RRT. It has a relatively large randomness, and in general, the narrower the feasible space, the larger $t_{\text{rrt}}$ tends to be. $t_{\text{sfc}}$ is the time it takes to generate an SFC using RILS based on the initial path, which is positively correlated with the number of path segments and the point cloud size. Since there are alternative methods available for the initial path search and the SFC generation that can be plug and play, we do not discuss the impact of these two stages on the efficiency of our framework here. $t_{\text{opt}}$ is the time it takes to solve the trajectory optimization problem, and $N_{\text{iter}}$ is the iteration number of L-BFGS. The time complexity of each L-BFGS iteration (where the value of the objective function and the related gradients are calculated) is $O(M)$ [12], so $t_{\text{opt}}$ is approximately proportional to $M$ and $N_{\text{iter}}$. The trajectory optimization stage shows relatively high solving efficiency, comparable to the performance on the CPU in [11], while [11] generates only 3-D position trajectories for underactuated MAVs. Therefore, our framework have the application potential in scenarios requiring real-time performance.
@@ IV-D Real-World Experiments
The results presented in this section aim to verify the applicability of the proposed method can be applied to real OMAVs. We set up cuboid obstacles with known sizes and positions in the environment. Then, we generate collision-free 6DoF trajectories in advance using the proposed method.
The experimental platform is the OmniHex (Fig. 5b) mentioned in Section II-C. Six Dynamixel XH430-W210-T servos each provide one rotor’s degree of tilt freedom. An OptiTrack motion capture system provides high-precision real-time 6-D pose feedback to the vehicle at 100Hz via WiFi. For the trajectory optimization settings, we set the kinodynamic constraints according to the real vehicle’s performance as : $v_{\max}=0.36\text{m}\cdot\text{s}^{-1},a_{\max}=0.5\text{m}\cdot\text{s}^{-2},\omega_{\max}=1.0\text{rad}\cdot\text{s}^{-1}$.
@@ V Conclusion
We present a computationally efficient 6-D trajectory generation framework which can fully exploit the obstacle avoidance potential of OMAVs. A 6-D trajectory optimization problem considering safety and kinodynamic constraints is formulated. A rational quaternion-based rotation parameterization method is adopted to achieve efficient optimization and high-quality solution. Simulations and real-world experiments are carried out to verify the performance of our method. Our method can be applied to any platform that can do free and controlled 6-D rigid body motion, including spacecraft. In the future, the goal is to build an OMAV with autonomous navigation ability in complex environments based on the proposed method, as well as the onboard sensing system. This will promote the application of OMAVs in aerial manipulators in cluttered environments.
Appendix A Gradient Calculation in Trajectory Optimization
Gradient of the time-regularized smoothness term $\mathcal{J}:=J+k_{\rho}\|\mathbf{T}\|_{1}$ can be calculated as follows:
$\frac{\partial\mathcal{J}}{\partial\mathbf{c}_{i}}=2\left(\int_{0}^{T_{i}}\bm{\beta}^{(s)}(t)\bm{\beta}^{(s)}(t)^{\top}\text{d}t\right)\mathbf{c}_{i},$ (28)
$\frac{\partial\mathcal{J}}{\partial T_{i}}=\mathbf{c}_{i}^{\top}\bm{\beta}^{(s)}(t)\bm{\beta}^{(s)}(t)^{\top}\mathbf{c}_{i}+k_{\rho}.$ (29)
For the penalty terms, each of them is the summary of several sub-penalty terms. When the constraint corresponding to a sub-penalty is satisfied, the sub-penalty and its gradient are always 0. So we only need to consider the sub-penalties that violate their corresponding constraints, and we denote them as
$P_{v_{ij}}:=\left(\|\dot{\mathbf{p}}(\hat{t}_{ij})\|_{2}^{2}-v_{\text{max}}^{2}\right)^{3}\frac{T_{i}}{\kappa}=\mathcal{G}_{v_{ij}}^{3}\frac{T_{i}}{\kappa},$ (30)
$P_{a_{ij}}:=\left(\|\ddot{\mathbf{p}}(\hat{t}_{ij})\|_{2}^{2}-a_{\text{max}}^{2}\right)^{3}\frac{T_{i}}{\kappa}=\mathcal{G}_{a_{ij}}^{3}\frac{T_{i}}{\kappa},$ (31)
$P_{\omega_{ij}}:=\left(\|\bm{\omega}(\hat{t}_{ij})\|_{2}^{2}-\omega_{\text{max}}^{2}\right)^{3}\frac{T_{i}}{\kappa}=\mathcal{G}_{\omega_{ij}}^{3}\frac{T_{i}}{\kappa},$ (32)
$P_{c_{ijlk}}:=\left(\mathbf{n}_{i,k}^{\top}\left(\mathbf{p}(\hat{t}_{ij})+\mathbf{R}(\hat{t}_{ij})\tilde{\bm{v}}_{l}\right)-d_{i,k}\right)^{3}\frac{T_{i}}{\kappa}=\mathcal{G}_{c_{ijlk}}^{3}\frac{T_{i}}{\kappa}.$ (33)
Obviously, the penalty terms corresponding to the $i$-th piece are only related to $\mathbf{c}_{i}$ and $T_{i}$. Then, the expressions are given as follows:
$\frac{\partial P_{v_{ij}}}{\partial\mathbf{c}_{i}}=\begin{bmatrix}\frac{6T_{i}}{\kappa}\mathcal{G}_{v_{ij}}^{2}\bm{\beta}^{(1)}(\frac{j}{\kappa}T)\dot{\mathbf{p}}^{\top}&\mathbf{0}_{2s\times 3}\end{bmatrix},$ (34a)
$\frac{\partial P_{v_{ij}}}{\partial T_{i}}=\frac{\mathcal{G}_{v_{ij}}^{2}}{\kappa}\left(\mathcal{G}_{v_{ij}}+\frac{6jT_{i}}{\kappa}\dot{\mathbf{p}}^{\top}\ddot{\mathbf{p}}\right),$ (34b)
$\frac{\partial P_{a_{ij}}}{\partial\mathbf{c}_{i}}=\begin{bmatrix}\frac{6T_{i}}{\kappa}\mathcal{G}_{a_{ij}}^{2}\bm{\beta}^{(2)}(\frac{j}{\kappa}T)\ddot{\mathbf{p}}^{\top}&\mathbf{0}_{2s\times 3}\end{bmatrix},$ (34c)
$\frac{\partial P_{a_{ij}}}{\partial T_{i}}=\frac{\mathcal{G}_{a_{ij}}^{2}}{\kappa}\left(\mathcal{G}_{a_{ij}}+\frac{6jT_{i}}{\kappa}\ddot{\mathbf{p}}^{\top}\dddot{\mathbf{p}}\right),$ (34d)
$\frac{\partial P_{\omega_{ij}}}{\partial\mathbf{c}_{i}}=\begin{bmatrix}\mathbf{0}_{2s\times 3}&\frac{6T_{i}}{\kappa}\mathcal{G}_{\omega_{ij}}^{2}\left(\omega_{x}\frac{\partial\omega_{x}}{\partial\mathbf{c}_{i}^{\sigma}}+\omega_{y}\frac{\partial\omega_{y}}{\partial\mathbf{c}_{i}^{\sigma}}+\omega_{z}\frac{\partial\omega_{z}}{\partial\mathbf{c}_{i}^{\sigma}}\right)\end{bmatrix},$ (34e)
$\frac{\partial P_{\omega_{ij}}}{\partial T_{i}}=\frac{\mathcal{G}_{\omega_{ij}}^{2}}{\kappa}\left(\mathcal{G}_{\omega_{ij}}+\frac{6T_{i}}{\kappa}\frac{\partial\bm{\omega}}{\partial T_{i}}\bm{\omega}\right),$ (34f)
$\frac{\partial P_{c_{ijlk}}}{\partial\mathbf{c}_{i}^{p}}=\frac{3T_{i}}{\kappa}\mathcal{G}_{c_{ijlk}}^{2}\bm{\beta}(\frac{j}{\kappa}T_{i})\mathbf{n}_{i,k}^{\top},$ (34g)
$\left[\frac{\partial P_{c_{ijlk}}}{\partial\mathbf{c}_{i}^{\sigma}}\right]_{m,n}=\frac{3T_{i}}{\kappa}\mathcal{G}_{c_{ijlk}}^{2}\mathbf{n}_{i,k}^{\top}\frac{\partial\mathbf{R}}{\partial[\mathbf{c}_{i}^{\sigma}]_{m,n}}\tilde{\bm{v}}_{l},$ (34h)
[…]
where $\mathbf{p}$, $\mathbf{R}$, $\bm{\omega}$, and their derivatives are taken as the values at $\hat{t}_{ij}$; $[\mathbf{F}]_{m,n}$ is the entry of matrix $\mathbf{F}$ with row index $m$ and column index $n$ and $m\in\{1,\cdots,2s\},n\in\{1,\cdots,6\}$.

### v2

@@ III-C Rotation Parameterization
where $\mathbf{U}:=\begin{bmatrix}-\mathbf{r}&w\mathbf{I}+\mathbf{r}^{\wedge}\end{bmatrix}\in\mathbb{R}^{3\times 4}$ and $\mathbf{G}:=\begin{bmatrix}\frac{\partial q_{w}}{\partial\bm{\sigma}}&\frac{\partial q_{x}}{\partial\bm{\sigma}}&\frac{\partial q_{y}}{\partial\bm{\sigma}}&\frac{\partial q_{z}}{\partial\bm{\sigma}}\end{bmatrix}\in\mathbb{R}^{3\times 4}$. Then, we can calculate the rotation-related penalty terms.
@@
Figure 7: Trajectory generation results. (a) shows a trajectory generated in Scenario B (from the top view), and (b) shows that in Scenario C (from the side view). We adjusted the transparency of the point cloud visualization in the bottom image of (b) so that the generated SFC and 6-D trajectory can be displayed more clearly.
Figure 8: The efficiency comparison between stereographic-projection-based (SP) method and Euler-angle-based (EA) method. The circles represent outliers outside the range $[Q_{1}-2\cdot IQR,Q_{3}+2\cdot IQR]$, where $Q_{1}$ is the lower quartile, $Q_{3}$ is the upper quartile, and $IQR=Q_{3}-Q_{1}$.
@@ IV-B Parameter Settings
To achieve whole-body obstacle avoidance, as Section III-A mentions, we approximate the vehicle’s shape as a convex polyhedron that envelops the body. Here we use a cuboid (Fig. 5a) whose dimensions are $l_{x}=l_{y}=1.1\text{m},l_{z}=0.42\text{m}$, according to the size of OmniHex. The coordinates of the eight vertices in $\mathcal{F}_{b}$ are $\tilde{\bm{v}}_{l}=\begin{bmatrix}\pm\frac{l_{x}}{2}&\pm\frac{l_{y}}{2}&\pm\frac{l_{z}}{2}\end{bmatrix}^{\top}$. Note that it is also possible to utilize a shape with closer proximity, such as a hexagon. However, the extent to which the solution space can be expanded using this type of shape rather than a cuboid is very limited, and more vertices may require additional computation. We set $s=4$ to ensure the smoothness of control inputs according to (8). The initial value of $\mathbf{q}^{\sigma}$ of trajectory optimization is set to $\mathbf{0}$. The other optimization settings for simulations are listed in Table I.
@@ IV-C Simulation Results
In this section, simulations are carried out in three virtual scenarios to test the proposed method’s effectiveness and efficiency. We generate a point cloud map for each scenario to provide obstacle information for our trajectory generation framework. The three virtual scenarios are described as follows.
Scenario A: We expect the vehicle to fly through a narrow, straight, vertical passage. The passage is 0.7m wide, which is narrower than the vehicle’s width ($l_{x}$ and $l_{y}$) and wider than its thickness ($l_{z}$).
@@
Scenario C: We expect the vehicle to pass through a $30\text{m}\times 10\text{m}\times 8\text{m}$ area containing 1000 randomly distributed floating cuboid obstacles.
@@
TABLE II: Efficiency Evaluation.
@@
C 19 24 38.1m 55ms 120ms 48 422ms
Note: $t_{\text{rrt}}$ is the time it takes to find an initial feasible path using RRT. $t_{\text{sfc}}$ is the time it takes to generate an SFC using RILS based on the initial path. $t_{\text{opt}}$ is the time it takes to solve the trajectory optimization problem, and $N_{\text{iter}}$ is the iteration number of L-BFGS.
Fig. 6a shows the 6-D trajectory generated in Scenario A and the corresponding kinodynamic properties under the above settings. The generated trajectory allows the vehicle to smoothly and appropriately change its attitude to adapt to the narrow space in the passage, avoiding collision while flying toward the target. Fig. 6b presents the constrained kinodynamic properties. Although the constraints are relaxed, they are effectively satisfied in the resulting trajectory. Moreover, the velocity norm reaches $v_{\text{max}}$ most of the time, which shows that our method can fully exploit the performance of OMAVs. Note that due to the softening step, the resulting trajectory may slightly exceed the constraints [10]. Nonetheless, in most cases the given constraints can be well satisfied and we can handle the possible violation by carefully tuning the optimizer parameters or by adding a certain margin when approximating the shape of the robot in practice.
@@
For efficiency evaluation, we list the computation times of the results in Fig. 6 and 7 in Table II. $t_{\text{rrt}}$ has a relatively large randomness, and in general, the narrower the feasible space, the larger $t_{\text{rrt}}$ tends to be. $t_{\text{sfc}}$ is positively correlated with the number of path segments and the point cloud size. Since there are alternative methods available for the initial path search and the SFC generation that can be plug and play, we do not discuss the impact of these two stages on the efficiency of our framework here. The time complexity of each iteration (where the value of the objective function and the related gradients are calculated) is $O(M)$ [12], so $t_{\text{opt}}$ is approximately proportional to $M$ and $N_{\text{iter}}$. The trajectory optimization stage shows relatively high solving efficiency, comparable to the performance on the CPU in [11], while [11] generates only 3-D position trajectories for underactuated MAVs. Therefore, our framework have the application potential in scenarios requiring real-time performance.
In order to demonstrate the efficiency of stereographic-projection-based rotation parameterization (SP method), we also implement the Euler-Angle-based method (EA method) for additional comparison. The comparison experiment is carried out in a virtual environment that is basically the same as Scenario C, except that we set three cases for the number of obstacles: 1000, 2000, and 3000. For each case, we repeat the following steps 100 times: generate a point cloud map and an SFC, and then utilize our trajectory optimizer to perform trajectory optimization within that SFC using both SP and EA methods under the same optimization settings. Therefore, the number of trajectory pieces obtained by the two methods is equal in each trial. The results are presented in Figure 8. From $t_{\text{opt}}$ we can see that the SP method exhibits significantly higher efficiency than the EA method with a more concentrated time distribution and much fewer iterations for convergence. In addition, an increase in the obstacle number will increase obstacle distribution density, leading to an increasing detour of the trajectories, which is consistent with the general increase in $M$ with obstacle number as shown in Figure 8. On the other hand, $t_{\text{opt}}/(M\cdot N_{\text{iter}})$ is generally stable, which is consistent with the linear complexity of the trajectory optimization algorithm. In general, compared with EA method, SP method has higher convergence speed and is more suitable for real-time applications.
@@ IV-D Real-World Experiments
The results presented in this section aim to verify the applicability of the proposed method can be applied to real OMAVs. We set up cuboid obstacles with known sizes and positions in the environment. Then, we generate collision-free 6-D trajectories in advance using the proposed method. The experimental platform is the OmniHex (Fig. 5b) mentioned in Section II-C. Six Dynamixel XH430-W210-T servos each provide one rotor’s degree of tilt freedom. An OptiTrack motion capture system provides high-precision real-time 6-D pose feedback to the vehicle at 100Hz via WiFi. For the trajectory optimization settings, we set the kinodynamic constraints according to the real vehicle’s performance as : $v_{\max}=0.36\text{m}\cdot\text{s}^{-1},a_{\max}=0.5\text{m}\cdot\text{s}^{-2},\omega_{\max}=1.0\text{rad}\cdot\text{s}^{-1}$.
@@ V Conclusion
We present a computationally efficient 6-D trajectory generation framework which can fully exploit the obstacle avoidance potential of OMAVs. A 6-D trajectory optimization problem considering safety and kinodynamic constraints is formulated. A rational quaternion-based rotation parameterization method is adopted to achieve efficient optimization and high-quality solution. Simulations and real-world experiments are carried out to verify the performance of our method. Furthermore, our method can be applied to any platform that can do free and controlled 6-D rigid body motion, including spacecraft.
The primary limitation of our method is its lack of consideration of possible dynamic changes of the environment, such as pedestrians, which limits its scope of application. In addition, utilizing the independence of each sampling point in the penalty function, the computational efficiency of our method can be further improved with parallel computation. Moreover, we hope to enable autonomous OMAV-based aerial manipulation based on the proposed method, as well as the onboard sensing system. Achieving the above improvements and goals will be our future work.